OpenStudy (anonymous):

4y + 1 = y + 22. What is the value of y?

OpenStudy (curry):

y=7

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

subtract the one, 4y=y+21 Now subtract the y, 3y=21 Divide by 3, y=7

OpenStudy (curry):

medal?

OpenStudy (anonymous):

But, 4y = 28. So y+22 would be 29?

OpenStudy (anonymous):

4y+1=y+22 4y+1=22+y-y 4y-y+1=22 3y+1-1=22-1 3y=21 3y/3=21/3 y=7

OpenStudy (anonymous):

ah doesnt matter lol